masc is 7 years older than sam and the sum their ages is 27. What are their ages?
Answer by Cromlix(4381)

(x + 7) + x = 27
x + 7 + x = 27
2x = 27 - 7
2x = 20
x = 10
Masc is 17yrs old
Sam is 10yrs old.
Hope this helps.
:-)